# Low-altitude Friendly-Jamming for Satellite-Maritime Communications via Generative AI-enabled Deep Reinforcement Learning

### v1

@@ 5.2.1 Advantages of SAC
General DRL approaches, such as the deep Q-network (DQN), are typically effective for discrete action space [57], whereas discretizing for continuous action space problems is not feasible. Moreover, trust region policy optimization (TRPO) can enhance policy stability by optimizing the trust region, while the computational complexity is higher [58]. In contrast, SAC is well-suited for continuous-time problems [59]. First, by incorporating maximum entropy theory, SAC promotes policy diversity, enabling better adaptation and exploration in complex environments and avoiding local optima. Second, the offline data update strategy further enhances sample efficiency through the iterative use of the replay buffer. Finally, SAC integrates the policy gradient approach with the Q-value function for policy updates, which improves both sample efficiency and training stability. Therefore, we select SAC as the framework for addressing the MDP.
@@ 5.2.2 Process of SAC
where $\log\pi(\boldsymbol{a}[t]|\boldsymbol{s}[t])$ denotes the entropy of the policy $\pi$, which promotes explorability. Moreover, $\alpha$ is the temperature parameter to balance the reward and entropy [60].
@@ 5.3.1 Motivation of TransSAC Algorithm
(i) Local Optimization for Long-Term Optimization Objectives: The formulated MDP with long-term optimization objectives requires global optimization by accounting for the decisions across multiple time slots and the cumulative effects. However, in conventional SAC, strategies and value networks often rely on the current state to determine actions, which may ignore the correlation between historical information and long-term objectives. Therefore, enhancing the capability of SAC to solve long-term optimization problems is necessary.
(ii) Preset Suboptimal Weights for Multiple Optimization Objectives: In the MDP, we aim to simultaneously optimize multiple objectives. Generally, conventional SAC uses fixed weights for each optimization objective, yet sensitive weight configurations may make it potentially difficult to adapt to the long-term dynamic problem. Moreover, improper weight configurations may lead to suboptimal performance. At this point, SAC may prioritize energy consumption over security, and vice versa. This makes it challenging to effectively balance the conflicting objectives when the preset weights are suboptimal.
@@ 13             Soft update the target state-value network;
5.3.2 Transformer-enhanced Learning Strategy
Different from step-by-step computations, the transformer, which is a GenAI approach, leverages the self-attention mechanism to capture all positions in the sequence [61][62]. In the long-term MDP, the transformer incorporates contextual information about past slots into the current input, enhancing the richness of current information. This approach allows the transformer to effectively preserve temporal information and optimize long-term objectives. Therefore, we use the transformer to process actions and states from the replay buffer, ensuring that the current decision aligns with long-term optimization objectives. Note that this approach does not conflict with the Markov property of the MDP, as the approach does not rely on complete historical information. Instead, the approach can enhance the representation of the current state and action, providing contextual information to optimize decision-making. The overall framework is presented in Algorithm 2, with the details are as follows.
@@ 5 Apply residual linking and layer normalization;
First, positional encoding is a key component used to give positional information for each element in the sequence [63]. In the MDP, positional encoding provides relative or absolute temporal positions for the inputs (states and actions), which can be denoted by [63]
$\mathcal{E}_{2i}=\sin(\frac{P[t]}{\varpi^{2i/d_{e}}}),$ (27a)
$\mathcal{E}_{2i+1}=\cos(\frac{P[t]}{\varpi^{2i/d_{e}}}),$ (27b)
Second, the self-attention mechanism is a key for the transformer, which can capture global information by calculating the similarity between each state and action and other states and actions. This allows algorithms to efficiently consider the subsequent impact of decisions, thereby optimizing the strategies of the long-term optimization problem. Accordingly, the self-attention mechanism is expressed by [63]
where $\boldsymbol{Q}=\boldsymbol{H}\boldsymbol{W}^{Q}$, $\boldsymbol{K}=\boldsymbol{H}\boldsymbol{W}^{K}$, and $\boldsymbol{V}=\boldsymbol{H}\boldsymbol{W}^{V}$ denote the query, key, and value matrices, respectively, with $\boldsymbol{W}^{Q},\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ indicate the corresponding learnable weight matrices, and $\boldsymbol{H}$ denotes the joint denotation of $\mathcal{S}$ and $\mathcal{A}$. Moreover, $\mathcal{U}_{softmax}$ is the normalization operation performed using the softmax function, and $d_{\boldsymbol{K}}$ denotes the dimension of the $\boldsymbol{K}$-matrix.
Third, the multi-head attention mechanism is used to compute multiple parallel attentions, which is defined by
Finally, the output after the self-attention layer is fed into a feed-forward neural (FFN), which performs an independent nonlinear transformation of the output at each time slot. As such, the FFN enhances the capability of the model to learn complex patterns, and the transformation is given by [63]
In summary, the transformer can effectively address long-term optimization problems by capturing relationships across distant time slots, thereby enhancing state and action representations for more effective policy generation.
@@ 5.3.3 Multi-armed bandit (MAB)-based Weight Optimization Scheme
Different from the preset weights, we employ an MAB-based weight optimization scheme that dynamically explores weights during the multi-objective optimization process. The MDP is dynamic with long-term optimization objectives, and balancing the conflicting objectives is crucial and challenging. Moreover, fixed and preset weights may prioritize one objective over others, causing algorithms to converge to a local optimum over time. In this case, the MAB-based scheme continuously explores diverse weights, allowing the algorithm to discover a broader solution space. Specifically, by using the $\varepsilon$-greedy strategy, MAB balances exploration and exploitation, reducing the risk of falling into local optima. In addition, MAB does not require complex models or enormous computational resources, enabling quick execution. This makes it well-suited for real-time optimization within the system. The general process is presented in Algorithm 3, and the details are introduced as follows.
@@
In summary, the MAB-based weight optimization scheme continuously explores diverse weights for MDP, effectively balancing the solutions among conflicting optimization objectives. This avoids inappropriate weights that could lead to suboptimal performance and maintains computational efficiency.
1 Initialize the reward $\mathcal{R}_{n}(a)$ for each arm, selection count $N(a)$ for each arm, and probability $\varepsilon$;
@@ 2 Randomly initialize the weights $\tau_{1}$ and $\tau_{2}$;
3 if  random $<\varepsilon$ then
@@ 5.4 Complexity Analysis of TransSAC Algorithm
  • Network Initialization: This process involves parameter initialization. The computational complexity is $\mathcal{O}(2|\theta|+|\Phi|)$, where $|\theta|$ and $|\Phi|$ are the number of parameters in the critic and actor networks, respectively [64].
@@
  • Transformer Executing: The complexity of executing the transformer is $\mathcal{O}(N_{t}M_{s}W_{t}D_{t})$, where $D_{t}$ indicates the data sampled from $\mathcal{D}$, and $W_{t}$ denotes the computational complexity of the transformer [65].
@@ 6.1.1 Parameter Settings
We conduct simulation experiments in Python 3.8 and Visual Studio Code 1.91 environments, and perform all the experiments on a server with AMD EPYC 7642 48-Core CPU, NVIDIA GeForce RTX 3090 GPU and 128 GB RAM.
In the simulation, we randomly initialize the UAV within an 80 m × 80 m feasible flight region, as the UAV could be on another mission before. Moreover, for TransSAC, each actor and critic network has two hidden layers and an output layer, with ReLU as the activation function, and the parameters are updated with the standard Adam optimizer. In addition, the batch size for sampling from the replay buffer is set to 128, and the number of attention heads and expansion multiplier of the hidden layer in the forward propagation are both set to 8. Additionally, the remaining primary parameters are presented in Table II [44][66].
@@
6.1.2 Baselines
To illustrate the performance of the TransSAC algorithm, a comparative approach and various comparison algorithms are as follows.
  • Non-UAV Approach: The approach relies on the LEO satellite sending signals to the legitimate vessel, without using UAV to interfere with the illegitimate vessel. As such, this comparison approach emphasizes the necessity of the UAV-assisted friendly-jamming approach in implementing secure LEO satellite-maritime communications.
  • State-of-the-art Comparison Algorithms: We select deep deterministic policy gradient (DDPG) [67], twin delayed deep deterministic policy gradient (TD3) [68], proximal policy optimization (PPO) [69], and conventional SAC as comparison algorithms. These algorithms are commonly used to solve dynamic optimization problems [64]. Specifically, DDPG leverages the advantages of policy gradient approaches and deep learning, using the actor critic structure to enhance policy learning, TD3 is a modification of the DDPG that improves performance through three aspects, including double Q-learning, delayed update, and target policy smoothing, and PPO improves the performance of the agent by optimizing the policy function and proposes trimming operations to maintain the stability and efficiency of training. The parameters of comparison algorithms are presented in Table III. Additionally, we set the total number of iterations in the aforementioned algorithms to $1\times 10^{6}$, and evaluate these algorithms every 80 iterations during the training process.
@@ 6.2.1 Hyper-parameters Results
Figure 5: Average reward of different hyper-parameters of the TransSAC algorithm. (a) LR. (b) Number of Neurons. (c) UR.
@@ 6.2.3 Comparison with Other Algorithms
Fig. 7 gives the optimization objective values obtained by different algorithms. Specifically, the TransSAC algorithm has an optimal average secrecy rate, achieving secure communications. Moreover, the average energy consumption of the UAV of the TransSAC is considerably lower, which is crucial in maritime environments where recharging UAVs is challenging. Therefore, the reliable security and minimal energy consumption make TransSAC a practical and effective algorithm. Notably, since TD3 is an improvement on the DDPG, the average energy consumption of the UAV of TD3 has a minimal enhancement in DDPG, resulting in similar values for both.
@@ 7 Conclusion
This work has studied secure LEO satellite-maritime communications by low-altitude friendly jamming. We have considered the system that utilizes UAV as the low-altitude platform to send jamming signals, mitigating the risk of satellite-transmitted data being eavesdropped by the illegitimate vessel. Considering the conflicting objectives, we have formulated the SSMCMOP to maximize the secrecy rate and minimize the energy consumption of the UAV simultaneously. To tackle the complex and dynamic problem, we have reformulated it into an MDP. Then, we have proposed the TransSAC algorithm, a GenAI-enable DRL approach that integrates transformer and MAB strategies to capture global dependencies and explore diverse weights. Simulation results have shown that the TransSAC outperforms other baselines, and achieves an optimal secrecy rate with minimal energy consumption of the UAV. Additionally, we have determined the suitable constraint values for the MDP. This work can be extended by incorporating the imperfect positions of eavesdropping vessels, which will be explored in future work.

### v2

@@ 5.2.1 Advantages of SAC
General DRL approaches, such as the deep Q-network (DQN), are typically effective for discrete action space [fan2020theoretical], whereas discretizing for continuous action space problems is not feasible. Moreover, trust region policy optimization (TRPO) can enhance policy stability by optimizing the trust region, while the computational complexity is higher [Schulman2015]. In contrast, SAC is well-suited for continuous-time problems [Zhang2021]. First, by incorporating maximum entropy theory, SAC promotes policy diversity, enabling better adaptation and exploration in complex environments and avoiding local optima. Second, the offline data update strategy further enhances sample efficiency through the iterative use of the replay buffer. Finally, SAC integrates the policy gradient approach with the Q-value function for policy updates, which improves both sample efficiency and training stability. Therefore, we select SAC as the framework for addressing the MDP.
@@ 5.2.2 Process of SAC
where $\log\pi(\boldsymbol{a}[t]|\boldsymbol{s}[t])$ denotes the entropy of the policy $\pi$, which promotes explorability. Moreover, $\alpha$ is the temperature parameter to balance the reward and entropy [Zhang2021].
@@ 5.3.1 Motivation of TransSAC Algorithm
(i) Strong Temporal Correlation: The vessels and LEO satellites are governed by constrained moving trajectories, and the UAV flight paths need to satisfy physical continuity constraints. These constraints make their current positions strictly limited by previous states, thereby causing the MDP to exhibit strong temporal correlation. At this point, conventional SAC primarily focuses on the current state and ignores the strong temporal correlation, which may result in ineffective decision-making. Note that this may cause policies to fall into local optima and further affect the performance of the optimization objectives.
(ii) Large-Scale State and Action Spaces: In the MDP, the UAV needs to process large state inputs and select action spaces. This results in the MDP having large-scale state and action spaces where multiple action dimensions need to be jointly optimized, resulting in combinatorial complexity. However, conventional SAC relies on trial-and-error learning to discover relevant patterns and optimize policies. As such, large-scale state and action spaces result in SAC requiring more training rounds to interact with the environment to discover effective policies, which reduces sampling efficiency.
(iii) Preset Suboptimal Weights for Multiple Optimization Objectives: In the MDP, we aim to simultaneously optimize multiple objectives. Generally, conventional SAC uses fixed weights for each optimization objective, yet sensitive weight configurations may make it potentially difficult to adapt to the long-term dynamic problem. Moreover, improper weight configurations may lead to suboptimal performance. At this point, SAC may prioritize energy consumption over security, and vice versa. This makes it challenging to effectively balance the conflicting objectives when the preset weights are suboptimal.
@@
5.3.2 Transformer-enhanced Learning Strategy
The transformer, a key GenAI approach, employs the self-attention mechanism to compute the global correlation weights across temporal samples, so that it can establish comprehensive feature interactions between multiple states [Han2021]. As such, by incorporating the transformer into standard DRL algorithms, we can enhance temporal modeling capabilities, allowing policy networks to explicitly capture long-term dependencies and sequential patterns. This process effectively addresses biases in locally correlated samples, thereby overcoming the challenge of local optima in conventional SAC. Furthermore, the transformer employs parallelized sequence processing to efficiently handle large-scale state and action spaces. Moreover, through multi-head attention mechanisms, the transformer can decompose high-dimensional actions and process different dimensions separately, improving the ability of the neural network to fit policy functions. Therefore, we use the transformer to process actions and states to improve the performance of the conventional SAC. Note that this approach does not conflict with the Markov property of the MDP, as the approach does not rely on complete historical information. Instead, the approach can enhance the representation of the current state and action, providing contextual information to optimize decision-making. The overall framework is presented in Algorithm 2, with the details as follows.
First, positional encoding is a key component used to give positional information for each element in the sequence [Vaswani2017]. In the MDP, positional encoding provides relative or absolute temporal positions for the inputs (states and actions), which can be denoted by [Vaswani2017]
$\mathcal{E}_{2i}=\sin(\frac{P[t]}{\varpi^{2i/d_{e}}}),$ (27a)
$\mathcal{E}_{2i+1}=\cos(\frac{P[t]}{\varpi^{2i/d_{e}}}),$ (27b)
Second, the self-attention mechanism is a key for the transformer, which can capture global information by quantifying the relevance between each action and current state. This allows algorithms to efficiently evaluate the subsequent impact of decisions, thereby solving the challenge of local optima caused by strong temporal correlation. Accordingly, the self-attention mechanism is expressed by [Vaswani2017]
where $\boldsymbol{Q}=\boldsymbol{H}\boldsymbol{W}^{Q}$, $\boldsymbol{K}=\boldsymbol{H}\boldsymbol{W}^{K}$, and $\boldsymbol{V}=\boldsymbol{H}\boldsymbol{W}^{V}$ denote the query, key, and value matrices, respectively, with $\boldsymbol{W}^{Q},\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ indicate the corresponding learnable weight matrices, and $\boldsymbol{H}$ denotes the joint denotation of $\mathcal{S}$ and $\mathcal{A}$. Moreover, $\mathcal{U}_{softmax}$ is the normalization operation performed using the softmax function, and $d_{\boldsymbol{K}}$ denotes the dimension of the $\boldsymbol{K}$-matrix. Note that the parallelized sequence processing capability of the self-attention mechanism can comprehensively analyze entire state-action sequences, effectively avoiding random and ineffective explorations and greatly improving sample efficiency in the large action space.
@@ 13      Soft update the target state-value network;
Third, multi-head attention mechanism decomposes high-dimensional actions into subspaces, enabling different attention heads to focus on different action dimensions. This approach can improve the fitting ability of the neural network, which is defined by
Finally, the output after the self-attention layer is fed into a feed-forward neural (FFN), which performs an independent nonlinear transformation of the output at each time slot. As such, the FFN enhances the capability of the model to learn complex patterns, and the transformation is given by [Vaswani2017]
In summary, the transformer-enhanced learning strategy can effectively address the challenge of strong temporal correlation through enhanced state and action representations while handling large-scale state and action spaces through parallelized processing. These capabilities make the transformer particularly well-suited for our large-scale and long-term optimization problem.
@@ 5.3.3 Multi-armed bandit (MAB)-based Weight Optimization Scheme
Different from the preset weights, we employ an MAB-based weight optimization scheme that dynamically explores weights during the multi-objective optimization process. The MDP is dynamic with long-term optimization objectives, and balancing the conflicting objectives is crucial and challenging. Moreover, fixed and preset weights may prioritize one objective over others, causing algorithms to converge to a local optimum over time. In this case, the MAB-based scheme continuously and diversely explores weights, allowing the algorithm to discover a broader solution space. Specifically, by using the $\varepsilon$-greedy strategy, MAB balances exploration and exploitation, reducing the risk of falling into local optima. In addition, MAB does not require complex models or enormous computational resources, enabling quick execution. This makes it well-suited for real-time optimization within the system. The general process is presented in Algorithm 3, and the details are introduced as follows.
@@
In summary, the MAB-based weight optimization scheme diversely explores weights for MDP, effectively balancing the solutions among conflicting optimization objectives. This avoids inappropriate weights that could lead to suboptimal performance and maintains computational efficiency.
1Initialize the reward $\mathcal{R}_{n}(a)$ for each arm, selection count $N(a)$ for each arm, and probability $\varepsilon$;
@@ 2 Randomly initialize the weights $\tau_{1}$ and $\tau_{2}$;
3 if random $\varepsilon$ then
@@ 5.4 Complexity Analysis of TransSAC Algorithm
  • Network Initialization: This process involves parameter initialization. The computational complexity is $\mathcal{O}(2|\theta|+|\Phi|)$, where $|\theta|$ and $|\Phi|$ are the number of parameters in the critic and actor networks, respectively [Zhang2024a].
@@
  • Transformer Executing: The complexity of executing the transformer is $\mathcal{O}(N_{t}M_{s}W_{t}D_{t})$, where $D_{t}$ indicates the data sampled from $\mathcal{D}$, and $W_{t}$ denotes the computational complexity of the transformer [Han2021].
@@ 6.1.1 Parameter Settings
We conduct simulation experiments in Python 3.8 and Visual Studio Code 1.91 environments, and perform all the experiments on a server with AMD EPYC 7642 48-Core CPU, NVIDIA GeForce RTX 3090 GPU, and 128 GB RAM.
In the simulation, we randomly initialize the UAV within an 80 m × 80 m feasible flight region, as the UAV could be on another mission before. Note that our simulations account for vessel trajectory variability through stochastic environmental influences and dynamic random seed strategies, while varying the initial positions of LEO satellites across training episodes. These enable the agent to explore diverse states, enhancing simulation realism and facilitating comprehensive performance evaluation. Moreover, for the TransSAC algorithm, each actor and critic network has two hidden layers and an output layer, with ReLU as the activation function, and the parameters are updated with the standard Adam optimizer. In addition, the batch size for sampling from the replay buffer is set to 128, and the number of attention heads and expansion multiplier of the hidden layer in the forward propagation are both set to 8. Additionally, the remaining primary parameters are presented in Table II [Deng2021][Li2023a].
@@
6.1.2 Baselines
To illustrate the performance of the TransSAC algorithm, we compare it with the theoretical optimal secrecy rate, a comparative approach, and various comparison algorithms as follows:
(a) LR.
(b) Number of Neurons.
(c) UR.
Figure 5: Average reward of different hyper-parameters of the TransSAC algorithm.
(i) Theoretical Optimal Secrecy Rate: Under the fixed LEO satellite orbit and vessel moving trajectories, we consider a theoretical scenario where legitimate vessels receive LEO satellite signals without UAV interference, and the eavesdropper receives the maximum jamming power from the UAV. This idealized scenario allows us to calculate a theoretical optimal secrecy rate. Note that due to the NP-hard complexity and temporal characteristics of the problem, it is nearly impossible to obtain the theoretical optimal secrecy rate within real-world time constraints. Thus, this comparison indicates that our approach can obtain a near-optimal secrecy rate, making it more suitable and valuable for the considered scenario.
(ii) Non-UAV Approach: The approach relies on the LEO satellite sending signals to the legitimate vessel, without using UAV to interfere with the illegitimate vessel. As such, this comparison approach emphasizes the necessity of the UAV-assisted friendly-jamming approach in implementing secure LEO satellite-maritime communications.
(iii) State-of-the-art Comparison Algorithms: We select deep deterministic policy gradient (DDPG) [Qiu2019], twin delayed deep deterministic policy gradient (TD3) [Fujimoto2018], proximal policy optimization (PPO) [Schulman2017], and conventional SAC as comparison algorithms. These algorithms are commonly used to solve dynamic optimization problems [Zhang2024a]. Specifically, DDPG leverages the advantages of policy gradient approaches and deep learning, using the actor-critic structure to enhance policy learning, TD3 is a modification of the DDPG that improves performance through three aspects, including double Q-learning, delayed update, and target policy smoothing, and PPO improves the performance of the agent by optimizing the policy function and proposes trimming operations to maintain the stability and efficiency of training. The parameters of comparison algorithms are presented in Table III. Additionally, we set the total number of iterations in the aforementioned algorithms to $1\times 10^{6}$, and evaluate these algorithms every 80 iterations during the training process.
@@ 6.2.1 Hyper-parameters Results
Figure 6: Secrecy rate obtained by our approach in comparison to the theoretical optimal secrecy rate.
@@
6.2.2 Comparisons with Theoretical Optimal Secrecy Rate
We compare the security performance of LEO satellite-maritime communications between our approach and a theoretical scenario. Fig. 6 illustrates the secrecy rate obtained by our approach in comparison to the theoretical optimal secrecy rate during an episode. It can be seen that our method approaches the theoretical optimal secrecy rate. Note that our approach solves a multi-objective optimization problem, which requires trade-offs between secrecy rate and energy consumption of the UAV. Thus, our approach achieves a near-optimal secrecy rate, offering a practical and valuable solution for the considered scenario.
@@ 6.2.4 Comparison with Other Algorithms
Fig. 8 gives the optimization objective values obtained by different algorithms. Specifically, the TransSAC algorithm has an optimal average secrecy rate, achieving secure communications. Moreover, the average energy consumption of the UAV of the TransSAC algorithm is considerably lower, which is crucial in maritime environments where recharging UAVs is challenging. Therefore, the reliable security and minimal energy consumption make TransSAC a practical and effective algorithm. Notably, since TD3 is an improvement on the DDPG, the average energy consumption of the UAV of TD3 has a minimal enhancement in DDPG, resulting in similar values for both.
@@ 6.2.5 Convergence Performance
7 Discussion
In this section, we further discuss the performance of the proposed approach, and the details are as follows:
  • The Extended Scenarios of Our Approach: Our approach can be extended to multiple UAV cases, achieving enhanced security performance and expanding maritime security coverage capabilities, and the details are presented in Appendix A.
  • Indicative Simulation Scenario: We present an indicative simulation scenario, including the trajectories of the LEO satellite, vessels, and the UAV, as well as the performance of key parameters, to enhance the clarity of our results. The detailed analysis and simulations are shown in Appendix B.
  • Fulfillment of Constraints of the SSMCMOP: We elaborate that our approach and various algorithms all satisfy the constraints of the SSMCMOP, and the detailed discussion is provided in Appendix C.
  • Implementation of Real-time Data Transmission of Our Approach: We demonstrate that our DRL-based framework with the offline training and online execution approach can optimize real-time transmission rates by adapting to changing environmental states. Moreover, we use a Raspberry Pi to confirm that our approach can achieve real-time data transmission. See Appendix D for details.
@@ 8 Conclusion
This work has studied secure LEO satellite-maritime communications by low-altitude UAV friendly-jamming. We have considered the system that utilizes a UAV as a low-altitude platform to send jamming signals, mitigating the risk of satellite-transmitted data being eavesdropped by the illegitimate vessel. Considering the conflicting objectives, we have formulated the SSMCMOP to maximize the secrecy rate and minimize the energy consumption of the UAV simultaneously. To tackle the dynamic and NP-hard problem, we have reformulated it into an MDP. Then, we have proposed the TransSAC algorithm, a GenAI-enabled DRL approach that integrates transformer and MAB strategies to capture long-term dependencies and diversely explore weights. Simulation results have shown that the TransSAC algorithm outperforms other comparative methods and algorithms, achieving maximum secrecy rate with minimal energy consumption of the UAV. Additionally, we have determined the suitable constraint values for the MDP. This work can be extended by incorporating the imperfect positions of eavesdropping vessels and leveraging a DRL-based framework to predict the position of the illegitimate user, which will be explored in future work.
[Uncaptioned image] Jiawei Huang received a BS degree in Software Engineering from Dalian Jiaotong University, and a MS degree in Software Engineering from Jilin University in 2019 and 2024, respectively. She is currently studying Computer Science at Jilin University to get a Ph.D. degree. Her current research interests are UAV networks and optimization.
[Uncaptioned image] Aimin Wang received Ph.D. degree in Communication and Information System from Jilin University, Changchun, China, in 2004. He is currently a professor at Jilin University. His research interests are wireless sensor networks and QoS for multimedia transmission.
[Uncaptioned image] Geng Sun (Senior Member, IEEE) received the B.S. degree in communication engineering from Dalian Polytechnic University, and the Ph.D. degree in computer science and technology from Jilin University, in 2011 and 2018, respectively. He was a Visiting Researcher with the School of Electrical and Computer Engineering, Georgia Institute of Technology, USA. He is a Professor in the College of Computer Science and Technology at Jilin University. Currently, he is working as a visiting scholar at the College of Computing and Data Science, Nanyang Technological University, Singapore. He has published over 100 high-quality papers, including IEEE TMC, IEEE JSAC, IEEE/ACM ToN, IEEE TWC, IEEE TCOM, IEEE TAP, IEEE IoT-J, IEEE TIM, IEEE INFOCOM, IEEE GLOBECOM, and IEEE ICC. He serves as the Associate Editors of IEEE Communications Surveys & Tutorials, IEEE Transactions on Communications, IEEE Transactions on Vehicular Technology, IEEE Transactions on Network Science and Engineering, IEEE Transactions on Network and Service Management and IEEE Networking Letters. He serves as the Lead Guest Editor of Special Issues for IEEE Transactions on Network Science and Engineering, IEEE Internet of Things Journal, IEEE Networking Letters. He also serves as the Guest Editor of Special Issues for IEEE Transactions on Services Computing, IEEE Communications Magazine, and IEEE Open Journal of the Communications Society. His research interests include Low-altitude Wireless Networks, UAV communications and Networking, Mobile Edge Computing (MEC), Intelligent Reflecting Surface (IRS), Generative AI and Agentic AI, and deep reinforcement learning.
[Uncaptioned image] Jiahui Li (Member, IEEE) received his B.S. in Software Engineering, and M.S. and Ph.D. in Computer Science and Technology from Jilin University, Changchun, China, in 2018, 2021, and 2024, respectively. He was a visiting Ph.D. student at the Singapore University of Technology and Design (SUTD). He currently serves as an assistant researcher in the College of Computer Science and Technology at Jilin University. His current research focuses on integrated air-ground networks, UAV networks, wireless energy transfer, and optimization.
[Uncaptioned image] Jiacheng Wang received the Ph.D. degree from the School of Communication and Information Engineering, Chongqing University of Posts and Telecommunications, Chongqing, China. He is currently a Research Associate in computer science and engineering with Nanyang Technological University, Singapore. His research interests include wireless sensing, semantic communications, and metaverse.
[Uncaptioned image] Dusit Niyato (M’09-SM’15-F’17) is a professor in the College of Computing and Data Science, at Nanyang Technological University, Singapore. He received B.Eng. from King Mongkuts Institute of Technology Ladkrabang (KMITL), Thailand and Ph.D. in Electrical and Computer Engineering from the University of Manitoba, Canada. His research interests are in the areas of mobile generative AI, edge general intelligence, quantum computing and networking, and incentive mechanism design.
[Uncaptioned image] Victor C. M. Leung (Life Fellow, IEEE) is currently a distinguished professor of computer science and software engineering with Shenzhen University, China. He is also an emeritus professor of electrial and computer engineering and the Director with the Laboratory for Wireless Networks and Mobile Systems, University of British Columbia. He has coauthored more than 1300 journal/conference papers and book chapters, and has been named in the current Clarivate Analytics list of Highly Cited Researchers. His research interests include the broad areas of wireless networks and mobile systems. Dr. Leung is also on the editorial boards of IEEE Transactions on Green Communications and Networking, IEEE Transactions on Cloud Computing, IEEE Access, and several other journals. He was the recipient of the IEEE Vancouver Section Centennial Award, 2011 UBC Killam Research Prize, 2017 Canadian Award for Telecommunications Research, 2018 IEEE TCGCC Distinguished Technical Achievement Recognition Award, and has coauthored papers that were the recipient of the 2017 IEEE ComSoc Fred W. Ellersick Prize, 2017 IEEE Systems Journal Best Paper Award, 2018 IEEE CSIM Best Journal Paper Award, and 2019 IEEE TCGCC Best Journal Paper Award. He is also the Life Fellow of IEEE, and a Fellow of the Royal Society of Canada, Canadian Academy of Engineering, and Engineering Institute of Canada.